# From Learning to Safety: A Direct Data-Driven Framework for Constrained Control

### v1

@@ 4.3 Learning SACBFs via RL
The optimal value function $B^{*}$, which can be obtained by dynamic programming, is a CBF with the safe set being the maximal control-invariant set [9]. Using Proposition 1 and letting $\beta_{B}=1$, we know that the state-action optimal value function of (4.3), defined by
$Q^{B*}(x,u):=\max\{h(x),\;B^{*}(f(x,u))\},$ (16)
$Q^{B}(x,u)=\max\{h(x),\min_{u^{+}\in U}Q^{B}(f(x,u),u^{+})\}.$ (17)
Proposition 3.
Consider the Bellman optimality equation (17). Any solution $Q^{B}$ satisfying $\min_{x\in\mathcal{X},u\in U}Q^{B}(x,u)\leq 0$ is an SACBF with the safe set $\mathcal{S}_{Q}=\{x\in\mathbb{R}^{n_{x}}|\min_{u\in U}Q^{B}(x,u)\leq 0\}$.
Proof.
It follows from $\min_{x\in\mathcal{X},u\in U}Q^{B}(x,u)\leq 0$ that $\mathcal{S}_{Q}$ is non-empty. From (17), we deduce that $h(x)\leq\min_{u\in U}Q^{B}(x,u)\leq 0,\;\forall x\in\mathcal{S}_{Q}$. The above arguments prove that $Q^{B}$ satisfies (i) of Definition 2. The second condition of Definition 2 directly follows from $\mathcal{S}_{Q}=\{x\in\mathbb{R}^{n_{x}}|\min_{u\in U}Q^{B}(x,u)\leq 0\}$. Finally, if $Q(x,u)\leq 0$, by (17) we have $\min_{u^{+}\in U}Q^{B}(f(x,u),u^{+})\leq 0$, which means that the successor state $f(x,u)\in\mathcal{S}_{Q}$. ∎
4.4 Special case when nonlinearity is known
The methods for synthesizing SACBFs proposed above are aimed at general nonlinear systems. When the nonlinearity of the system is known, we can leverage formulas from data-driven control [31, 32] and compute SACBFs through convex optimization. Formally, in this subsection, we require the system to satisfy the following conditions:
Assumption 3.
  • The system is described by the following equation:
    $x_{t+1}=A\bar{v}(x_{t})+Bu_{t},$ (19)
    where $\bar{v}(x)=\left[\begin{array}[]{c}x\\ v(x)\end{array}\right]$ includes a known nonlinear function $v(\cdot):\mathbb{R}^{n_{x}}\to\mathbb{R}^{n_{v}}$. Besides, the linear system obtained by linearizing (19) at the origin is controllable in the absence of the constraints.
  • The constraints are linear, i.e., $X=\{x\in\mathbb{R}^{n_{x}}|H_{x}x\leq h_{x}\},\;U=\{u\in\mathbb{R}^{n_{u}}|H_{u}u\leq h_{u}\}$. Here, $H_{x}$ and $H_{u}$ are matrices with appropriate dimensions. Besides, $h_{x}$ and $h_{u}$ are two vectors with strictly positive elements. This indicates that the origin is contained in the interior of $X\times U$.
  • The target region $X_{\mathrm{tar}}=\{0\}$.
To begin with, we recall the definition of Hankel matrix [31] associated to any time-varying signal $\{z_{t}\}_{t\in\mathbb{N}}$ with $z_{t}\in\mathbb{R}^{n_{z}},\;t\in\mathbb{N}$:
$\mathbf{z}_{i,j,k}=\left[\begin{array}[]{cccc}z_{i}&z_{i+1}&\cdots&z_{i+k-j}\\ z_{i+1}&z_{i+2}&\cdots&z_{i+k-j+1}\\ \vdots&\vdots&\ddots&\vdots\\ z_{j+i-1}&z_{j-i}&\cdots&z_{k+i-1}\end{array}\right].$ (20)
Then, the Hankel matrices $\mathbf{x}_{i,j,k}$ and $\mathbf{u}_{i,j,k}$ contain the state and input data collected from a simulation of the system in (19). The following lemma shows that (19) has equivalent data-based representations.
Lemma 2 ([32]).
Let $\mathbf{D}:=\left[\begin{array}[]{l}\mathbf{u}_{0,1,k}\\ {\mathbf{\bar{v}}_{0,1,k}}\end{array}\right]=\left[\begin{array}[]{l}\mathbf{u}_{0,1,k}\\ \mathbf{x}_{0,1,k}\\ \mathbf{v}_{0,1,k}\end{array}\right]$. Suppose that $\mathrm{rank}(\mathbf{D})=n_{x}+n_{u}+n_{v}$ and that Assumption 3 holds. Then, the system in (19) has the following equivalent representation:
$x_{t+1}=\mathbf{x}_{1,1,k}\mathbf{D}^{\dagger}\left[\begin{array}[]{l}u_{t}\\ \bar{v}(x_{t})\end{array}\right].$ (21)
Furthermore, the system in closed loop with $u=K\bar{v}(x)$ has the following equivalent representation:
$x_{t+1}=\mathbf{x}_{1,1,k}G\bar{v}(x_{t})$ (22)
with $G=[G_{1}\;G_{2}]$, $G_{1}\in\mathbb{R}^{k\times n_{x}}$, and $G_{2}\in\mathbb{R}^{k\times n_{v}}$satisfying
$\left[\begin{array}[]{l}K\\ I_{n_{x}+n_{v}}\end{array}\right]=\mathbf{D}G.$ (23)
In Lemma 2, the rank condition on $\mathbf{D}$ is referred to as the condition of persistency of excitation, which is commonly assumed among direct data-driven control approaches [31]. This condition implies that the collected data is sufficiently informative to represent the system behavior. Next, we show that under persistency of excitation and a mild assumption on the nonlinear term $v$ (See Assumption 4 below), we can formulate a data-driven computation of an SACBF in terms of a convex optimization problem subject to LMI constraints.
Assumption 4.
There exist a constant $\eta>0$ and a matrix $M\in\mathbb{R}^{n_{x}\times n_{x}}$ such that $v^{T}(x)v(x)\leq\eta x^{T}M^{T}Mx$ $\forall x\in X$.
This assumption limits the magnitude of the nonlinear term $v(x)$ by a quadratic function $\eta x^{T}M^{T}Mx$. Such an assumption is commonly encountered in the literature addressing the stabilization of nonlinear systems using LMIs [33].
Consider
$Q^{B}_{P}(x,u):=$
$\max\left\{h(x),[u^{T}\;\bar{v}^{T}(x)]\mathbf{D}^{\dagger T}\mathbf{x}^{T}_{1,1,k}P^{\!-\!1}\mathbf{x}_{1,1,k}\mathbf{D}^{\dagger}\left[\begin{gathered}u\\ \bar{v}(x)\end{gathered}\right]\!-\!1\right\},$ (26)
where $P\in\mathbb{R}^{n_{x}\times n_{x}}$ is a positive-definite matrix. The consideration of the parameterization (4.4) is inspired from Proposition 1 and (21), in which we specify $h(x)=\max_{i}\{H_{x,i}x-h_{x,i}\}$ so that $\{x\in\mathbb{R}^{n_{x}}|h(x)\leq 0\}$ is identical to $\{x\in\mathbb{R}^{n_{x}}|H_{x}x\leq h_{x}\}$. The second argument of the max function indicates that we firstly consider a quadratic CBF candidate $x^{T}P^{-1}x-1$, and then use the open-loop representation (21) and the result in Proposition 1 to get an SACBF candidate.
Now, we consider the following convex optimization problem:
$\min_{P,R_{1},G_{2}}$ $\;-\log\operatorname{det}(P)$ (27a)
$\mathrm{s.t.}\;$ $\left[\begin{array}[]{cccc}P&0&R^{T}_{1}\mathbf{x}^{T}_{1,1,k}&PM^{T}\\ &I_{n_{v}}&G^{T}_{2}\mathbf{x}^{T}_{1,1,k}&0\\ &*&P&0\\ &*&*&I_{n_{x}}/\eta\end{array}\right]\succeq 0$ (27f)
$\left[\begin{array}[]{cccc}h^{2}_{u,i}P&0&R^{T}_{1}\mathbf{u}^{T}_{0,1,k}H^{T}_{u,i}&PM^{T}\\ &I_{n_{v}}&G^{T}_{2}\mathbf{u}^{T}_{0,1,k}H^{T}_{u,i}&0\\ &*&1&0\\ &*&*&I_{n_{x}}/\eta\end{array}\right]\succeq 0$ (27k)
$\quad\quad i=1,...,\dim(h_{u})$ (27l)
$\left[\begin{array}[]{cc}h^{2}_{x,i}&H_{x,i}P\\ &P\end{array}\right]\succeq 0,\;i=1,...,\dim(h_{x})$ (27o)
$\left[\begin{array}[]{c}\mathbf{x}_{0,1,k}\\ \mathbf{v}_{0,1,k}\end{array}\right]R_{1}=\left[\begin{array}[]{c}P\\ 0\end{array}\right]$ (27t)
$\left[\begin{array}[]{c}\mathbf{x}_{0,1,k}\\ \mathbf{v}_{0,1,k}\end{array}\right]G_{2}=\left[\begin{array}[]{c}0\\ I_{n_{v}}\end{array}\right]$ (27y)
$P\succeq 0,\;R_{1}\in\mathbb{R}^{k\times n_{x}},\;G_{2}\in\mathbb{R}^{k\times n_{v}},$ (27z)
where “*” indicates that the lower triangular part of the matrix is the symmetric counterpart of the upper triangular part and $H_{x,i},\;h_{x,i}$ refer to the $i$th row of $H_{x},\;h_{x}$. Based on the optimization problem (27), we can construct an SACBF $Q^{B}_{P}$, as stated in the next result.
Proposition 4.
Consider the system (19), the SACBF parameterization (4.4), and the optimization problem (27). Suppose that $\mathrm{rank}(\mathbf{D})=n_{x}+n_{u}+n_{v}$ and Assumptions 3-4 hold. Then, problem (27) is feasible, and its optimal solution $P^{*}$ makes $Q^{B}_{P^{*}}$ an SACBF with the corresponding safe set $\{x\in\mathbb{R}^{n_{x}}|x^{T}P^{*-1}x\leq 1\}$.
Proof.
The proof is based on the standard argument of using LMIs to synthesize CBFs [23]. By letting $R_{1}=G_{1}P$ and $K=\mathbf{u}_{0,1,k}G$, (27t) and (27y) are equivalent to (23).
We consider the condition
$x^{+T}P^{-1}x^{+}\leq 1,\;\forall x^{T}P^{-1}x\leq 1,\;\forall v^{T}(x)v(x)\leq\eta x^{T}M^{T}Mx,$ (28)
where $x^{+}$ is the successor state of the closed-loop system (22). By sequentially using the S-procedure [34], utilizing the Schur complement [34], multiplying both sides by $\mathrm{diag}([P\;I_{n_{v}}\;I_{n_{x}}])$, and applying the Schur complement once more, we get that (27f) is sufficient to ensure (28). The condition (28) further implies the invariance of the safe set for the closed-loop system (22).
Then, by a similar argument we can prove that (27l) is a sufficient condition for
$H_{u}K\bar{v}(x)\leq h_{u},\;\forall x\in\{x\in\mathbb{R}^{n_{x}}|x^{T}P^{-1}x\leq 1\}$ (29)
to hold. The condition (29) means that the controller $u=K\bar{v}(x)$ satisfies the input constraint for all states in the safe set. Besides, (27o) is equivalent to the condition
$\{x\in\mathbb{R}^{n_{x}}|x^{T}P^{-1}x\leq 1\}\subseteq X.$ (30)
The above arguments prove that $B:=x^{T}P^{*-1}x-1$ is a CBF. Meanwhile, by comparing the right-side of (4.4) and the form of $B$, and noting that the open loop system (19) can be represented by (21), we have the relation $Q^{B}_{P^{*}}(x,u)=\max\{h(x),\;B(f(x,u))$. Finally, according to Proposition 1, we can conclude that $Q^{B}_{P^{*}}$ is an SACBF.
We now prove the feasibility of (27). The feasibility of (27t) and (27y) follows from (23). Regarding the other constraints, since (i) the linearized system of (19) is controllable, (ii) $\lim_{x\rightarrow 0}\frac{||v(x)||_{2}}{||x||_{2}}=0$ from Assumption 4, and (iii) the origin is within the interior of $X$, we can always replace $P$ by $\alpha_{P}P$ with $\alpha_{P}\in(0,1)$ to make the conditions (28)-(30) satisfied. This proves the feasibility of (27f)-(27o). ∎
@@ 5 Performance analysis and guarantee under learning errors
$Q^{B}_{\omega}(x,\pi_{\mathrm{s}}(x))\leq\beta q_{\omega}(x)+\varepsilon,\;\forall x\in\mathcal{S}_{0}$ (33a)
$q_{\omega}(f(x,u))\leq Q^{B}_{\omega}(x,u)+\varepsilon,\;\forall(x,u)\in\mathcal{S}_{0}\times U$ (33b)
$q_{\omega}(x)\geq\frac{2}{1-\beta}\varepsilon,\;\forall x\in\mathcal{X}\setminus\mathcal{S}_{0}$ (33c)
@@
(33c) leads to $\mathcal{S}_{\omega}\subseteq\mathcal{S}_{0}$. Consider any $x\in\mathcal{S}_{\omega}$, we have
@@
The condition in (32) quantifies the local approximation quality of the regression model $Q^{B}_{\omega}$, which is often assumed in SL literature [36, 37, 23]. The two conditions in (33a) and (33b) are motivated by (9c) and (9d), where we assume that the constraints in (9c) and (9d) are violated and the degree of violation is limited by $\varepsilon$. Besides, the condition (33c) is also justifiable, as it is introduced to ensure that the safe set of $Q^{B}_{\omega}$ is a subset of $\mathcal{S}_{0}$.
@@
Different from existing work that primarily addresses the robustness and ISSf for learning-based control under external disturbances [5], input disturbances [37], or inaccurate observations [38], our findings offer a systematic way for designers to guarantee the safety of the relaxed safety-filtered policy $\pi^{\mathrm{r}}_{\theta,\omega}$ when there are perturbations on the safety constraints, by appropriately tightening the state constraint. More importantly, we explicitly quantify the degrees of tightening and relaxation, based on the parameters $\beta$ and $\beta_{B}$ of the given CBF and the safe policy $\pi_{\mathrm{s}}$ as well as the bound $\varepsilon$ on the learning error. In particular, to ensure safety for the SL method, the state constraint must be tightened to $h(x)+\frac{2\beta_{B}}{1-\beta_{B}}\varepsilon\leq 0$, while the safety constraint in the safety filter should be relaxed by adjusting $Q^{B}_{\omega}(x,u)\leftarrow Q^{B}_{\omega}(x,u)-\frac{1+\beta_{B}}{1-\beta_{B}}\varepsilon$. Similarly, in the expert-guided learning method, the state constraint should be tightened to $h(x)+\frac{2}{1-\beta}\varepsilon\leq 0$, and (9b) should be adjusted to (33c), with the relaxed SACBF $Q^{B}_{\omega}(x,u)\leftarrow Q^{B}_{\omega}(x,u)-\frac{1+\beta}{1-\beta}\varepsilon$ in the safety filter.
@@
The proposed constraint tightening approach is inspired by robust MPC [37], which ensures both recursive feasibility and safety. In contrast, our optimization-based control framework enforces safety through an SACBF constraint, rather than imposing state constraints over a finite prediction horizon. As a result, recursive feasibility alone does not guarantee safety, highlighting the need for constraint relaxation. Broadly speaking, state constraint tightening contributes to safety, whereas SACBF constraint relaxation addresses feasibility concerns.
@@ 6 Refining policies with SACBF constraints
A convenient approach for designing $Q_{\theta}$, allowing $\theta$ to be updated using most existing unconstrained policy-based RL algorithms is to consider a Euclidean distance objective function $Q_{\theta}(x,u):=||u-\pi_{\theta}(x)||_{2}$, where $\pi_{\theta}:\mathcal{X}\rightarrow\mathcal{U}$ is an explicit controller (e.g., a neural network controller) with the parameter $\theta$ [7]. This approach, however, could significantly compromise the optimality of the projected policy $\pi_{\theta,\omega}$. A less conservative approach is to make $Q_{\theta}$ approximate the constrained optimal value function $Q^{*}$ defined in (6), so that the policy derived from (3) more closely approximates the optimal constrained policy. To this end, we present a unified approach to transform unconstrained value-based RL algorithms into constrained ones, utilizing the obtained SACBF.
Normally, to obtain $Q^{*}$, as is shown in [40, Chapter 7.4.1], one typically needs to recursively compute the backward reachable sets $X_{k},\;k=0,1,...$ and enforce the constraint $f(x,u)\in X_{k}$ during the Bellman iteration. This procedure is in general intractable for nonlinear systems. Instead, we replace the state constraints $h(x_{t})\leq 0,\;\forall t\in\mathbb{N}$ in (3) with the obtained SACBF constraints $Q_{\omega}^{B}(x_{t},\pi(x_{t}))\leq 0,\;\forall t\in\mathbb{N}$. Note that this will be an equivalent transformation if $Q_{\omega}^{B}=Q^{B*}$. The benefit of this transformation is that the constraints during the Bellman iteration become fixed to $Q_{\omega}^{B}(x_{t},\pi(x_{t}))\leq 0$. This is a result of the inherent invariance of the safe set of $Q_{\omega}^{B}$.
@@
where $\Gamma$ is the Bellman operator. It is easy to show that $\Gamma$ is a monotonous contraction mapping. As a consequence, the uniqueness of the solution to (6) holds and the constrained Bellman iteration $\bar{Q}_{k+1}=\Gamma\bar{Q}_{k}$ converges to $\bar{Q}^{*}$ for any real-valued and bounded $\bar{Q}_{0}:\mathcal{X}\times\mathcal{U}\to\mathbb{R}$ [39, Proposition 2.1.1].
@@
where $\text{where }q_{\theta}(x,u)=g(x,u)+\gamma$ $\min_{u^{+}\in U}Q_{\theta}\left(f(x,u),u^{+}\right)$
[…]
@@
7 Comparison with other approaches and limitations
Comparison with learning CBFs: Machine learning has recently been developed for synthesizing CBFs (safety filters) with or without model uncertainties. A common methodology to synthesizing permissive CBFs is to link CBFs with value functions of optimal control problems, such as the approaches in [8, 12, 5, 17] and the current paper. The value function can then be approximated by RL or SL. Some authors assume the knowledge of a nominal model and a valid CBF for that model [14, 13]. Then, RL is used to learn the discrepancies between the CBF constraint associated with the nominal model and that related to the real model. However, all the above methods can only enable the design of model-based safety filters. This limitation is due to the inherent nature of CBFs, which need the model information to enforce the invariance condition. In contrast, in the current paper we fundamentally propose a new direct data-driven safe control framework, in which an SACBF is employed to evaluate the feasibility of input signals without using model information.

### v2

@@ 4.3 Supervised learning SACBFs from CBFs
The SL method is intended for scenarios where a valid CBF is already available, but the model is unknown. In such a case, even when a CBF is known, traditional safety filter design based on CBFs still depends on a nominal model (typically obtained through system identification) to enforce the invariance condition [48]. In contrast, the proposed SL method offers an alternative by learning a scalar SACBF to enforce invariance directly, avoiding the need for system identification.
@@ 5 Performance analysis and guarantee under learning errors
From Definitions 2 and 3, it is clear that if $\hat{Q}^{B}$ is an $\varepsilon$-ESSf SACBF, then $\hat{Q}^{B}-\kappa(\varepsilon)$ is an SACBF.
@@
$Q^{B}_{\omega}(x,\pi_{\mathrm{s}}(x))\leq\beta q_{\omega}(x)+\varepsilon,\;\forall x\in\mathcal{S}_{\omega}$ (22a)
$q_{\omega}(f(x,u))\leq Q^{B}_{\omega}(x,u)+\varepsilon,\;\forall(x,u)\in\mathcal{S}_{\omega}\times U$ (22b)
${\color[rgb]{0,0,1}\min_{x\in\mathcal{X}}q_{\omega}(x)\leq\frac{2}{1-\beta}\varepsilon}$ (22c)
@@
Consider any $x\in\mathcal{S}_{\omega}$, we have
@@
Theorems 1 and 2 provide a systematic way to explicitly quantify how much the degrees of constraint tightening and relaxation are sufficient to ensure safety for the controller $\pi^{\mathrm{r}}_{\theta,\omega}$ obtained from the relaxed safety filter (5). In particular, Theorem 1 states that if the approximation $Q^{B}_{\omega}$ satisfies condition (21), to ensure safety for the SL method, the state constraint must be tightened to $h(x)+\frac{2\beta_{B}}{1-\beta_{B}}\varepsilon\leq 0$, while the safety constraint in the safety filter should be relaxed to $Q^{B}_{\omega}(x,u)\leq\frac{1+\beta_{B}}{1-\beta_{B}}\varepsilon$. Similarly, in the expert-guided learning method, Theorem 2 tells that the state constraint should be tightened to $h(x)+\frac{2}{1-\beta}\varepsilon\leq 0$, with the relaxed SACBF constraint $\leftarrow Q^{B}_{\omega}(x,u)\leq\frac{1+\beta}{1-\beta}\varepsilon$ in the safety filter.
The safety performance analysis in Theorems 1 and 2 is only valid when $\beta$ is bounded substantially below 1. Otherwise, the tightened state constraint $X_{\omega}$ shrinks drastically and can even become empty. However, for the expert-guided learning approach, this limitation can be mitigated in practice since $\beta$ can be tuned through an outer-loop procedure to maintain a value sufficiently below 1.
@@
The soundness of Theorems 1 and 2 relies on conditions (21) and (22). Condition (21) quantifies the local approximation quality of the regression model, which is often assumed in SL literature [37, 25, 24]. The two conditions in (22a) and (22b) are motivated by (14c) and (14d), where we assume that the constraints in (14c) and (14d) are violated, and the degree of violation is limited by $\varepsilon$. Besides, condition (22c) is also justifiable, as it is introduced to ensure that $\mathcal{S}_{\omega}$ is non-empty. Condition (22c) can be easily checked by solving a nonlinear optimization problem, while checking the other conditions and computing $\varepsilon$ can be achieved by using existing deterministic or probabilistic verification tools [25, 23, 44].
@@
Classical ISSf [28] or MPC-based tightening techniques [25] operate in the state space and rely on explicit knowledge of system dynamics. In contrast, our ESSf framework is built upon SACBFs, which evaluate safety directly on $(x,u)$ pairs. This fundamental shift eliminates the need for model predictions and allows safety analysis entirely from data. To the best of our knowledge, such an error-to-state formulation for state-action certificates in a direct data-driven setting has not appeared in prior work.
Compared to robust MPC [25], which typically relies on constraint tightening, ensuring both safety and recursive feasibility for our optimization-based controller is more challenging. In MPC, recursive feasibility directly guarantees constraint satisfaction of the resulting policy, but this implication does not generally hold for the proposed controller. To address this challenge, we introduce a novel constraint tightening-relaxation approach. Our analysis provides closed-form, quantitative relationships between learning errors in the SACBF, the required tightening of state constraints, and the necessary relaxation of the SACBF constraint.
@@ 6 Refining policies with SACBF constraints
[…]
Normally, to obtain $Q^{*}$, as is shown in [5, Chapter 7.4.1], one typically needs to recursively compute the backward reachable sets $X_{k},\;k=0,1,$... and enforce the constraint $f(x,u)\in X_{k}$ during the Bellman iteration. This procedure is in general intractable for nonlinear systems. Instead, we replace the state constraints $h(x_{t})\leq 0,\;\forall t\in\mathbb{N}$ in (3) with the obtained SACBF constraints $Q_{\omega}^{B}(x_{t},\pi(x_{t}))\leq 0,\;\forall t\in\mathbb{N}$. Note that this will be an equivalent transformation if $Q_{\omega}^{B}=Q^{B*}$. The benefit of this transformation is that the constraints during the Bellman iteration become fixed to $Q_{\omega}^{B}(x_{t},\pi(x_{t}))\leq 0$. This is a result of the inherent invariance of the safe set of $Q_{\omega}^{B}$.
@@
where $\Gamma$ is the Bellman operator. It is easy to show that $\Gamma$ is a monotonous contraction mapping. As a consequence, the uniqueness of the solution to (6) holds and the constrained Bellman iteration $\bar{Q}_{k+1}=\Gamma\bar{Q}_{k}$ converges to $\bar{Q}^{*}$ for any real-valued and bounded $\bar{Q}_{0}:\mathcal{X}\times\mathcal{U}\to\mathbb{R}$ [4, Proposition 2.1.1].
@@
$\text{where }q_{\theta}(x,u)=g(x,u)+\gamma$ $\min_{u^{+}\in U}Q_{\theta}\left(f(x,u),u^{+}\right)$
$\text{s.t. }Q_{\omega}^{B}\left(f(x,u),u^{+}\right)\leq 0.$ (25b)
@@
Under the additional assumption that for any $k\in\mathbb{N}$, the deviation $|Q_{\theta_{k+1}}(x,u)-\Gamma Q_{\theta_{k}}(x,u)|$ is upper-bounded, a standard condition in the literature on RL convergence analysis (e.g., [4, Section 2.3.1]), it follows that $Q_{\theta_{k}}$ converges to a neighborhood of $\bar{Q}^{*}$ as $k\to\infty$ [4, Proposition 2.3.2]. Moreover, note that the update of $Q_{\theta}$ does not affect the guarantees of safety or recursive feasibility of (5).
7 Discussions
Computational complexity: In our work, we use two deep neural networks $Q_{\theta}$ and $Q^{B}_{\omega}$ to parameterize the objective function $Q$ and the constraint function $Q^{B}$ in (7), respectively. Let $Q_{\theta}$ have $L_{Q}$ layers with widths $(d_{0},\dots,d_{L_{Q}})$, where $d_{0}=n_{x}+n_{u}$ and $d_{L_{Q}}=1$, and analogously $Q^{B}$ has widths $(\tilde{d}_{0},\dots,\tilde{d}_{L_{B}})$. The number of parameters of the two networks is
$P_{Q}=\sum_{l=1}^{L_{Q}}(d_{l-1}d_{l}+d_{l}),\qquad P_{B}=\sum_{l=1}^{L_{B}}(\tilde{d}_{l-1}\tilde{d}_{l}+\tilde{d}_{l}).$
A forward pass performs (approximately) one multiply and one add per weight, so the number of floating-point operations (FLOPs) per evaluation satisfies
$C_{\mathrm{for}}=\mathcal{O}(P_{Q}+P_{B}).$
Now assume that we solve the nonlinear program in the optimization-based controller by an iterative first-order method (e.g., projected gradient, penalty or barrier method). Let $K$ denote the number of iterations per control step. Each iteration requires one forward and backward pass through both networks, and the cost of a backward pass is on the same order as the forward pass, i.e., $C_{\mathrm{back}}=C_{\mathrm{for}}$. In addition, some cheap vector operations are needed to handle the constraint $u\in U$ (e.g., projections, line search updates), which takes $\mathcal{O}(n_{u})$ FLOPs. Usually, $P_{Q}+P_{B}\gg n_{u}$. Hence, over $K$ iterations, the total complexity per control step is
$C_{\mathrm{comp}}=\mathcal{O}\bigl(K(P_{Q}+P_{B})\bigr).$
Memory usage: We distinguish static (always present) and dynamic (per evaluation/iteration) memory.
The static memory is the memory used for storing the networks:
$M_{\mathrm{sta}}=\mathcal{O}(P_{Q}+P_{B}).$
During a forward and backward pass, activations, gradients, and the final solution (control input) must be stored. Hence the total dynamic memory per iteration is
$M_{\mathrm{dyn}}=\mathcal{O}\left(\sum_{l=1}^{L_{Q}}d_{l}+\sum_{l=1}^{L_{B}}\tilde{d}_{l}+n_{u}\right),$
Since $P_{Q}+P_{B}$ dominates $\sum_{l=1}^{L_{Q}}d_{l}+\sum_{l=1}^{L_{B}}\tilde{d}_{l}+n_{u}$, we can summarize that the total memory is
$M_{\mathrm{total}}=\mathcal{O}(P_{Q}+P_{B}).$
Comparison with learning CBFs: Machine learning has recently been developed for synthesizing CBFs (safety filters) with or without model uncertainties. A common methodology to synthesizing permissive CBFs is to link CBFs with value functions of optimal control problems, such as the approaches in [17, 9, 16, 46] and the current paper, or value functions of known policies [45, 55]. Some authors assume the knowledge of a nominal model and a valid CBF for that model [47, 53]. Then, RL [47, 53] or Gaussian process learning [8] is used to learn the discrepancies between the CBF constraint associated with the nominal model and the real model. By leveraging the covariance of the estimation provided by the Gaussian process and developing an event-trigger mechanism, [8] ensures the recursive feasibility of the safety filter in high probability [8]. Another approach is to formulate the synthesis of CBFs as a robust optimization problem and to use sampling-based methods to solve it [44, 38]. This approach is more sample-efficient than RL. However, all the above methods can only enable the design of model-based safety filters. In contrast, in the current paper we fundamentally propose a new direct data-driven safe control framework, in which an SACBF is employed to evaluate the feasibility of input signals without using model information.